# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v2

@@ Abstract.
In a hyper-relational knowledge graph (HKG), each fact is composed of a main triple associated with attribute-value qualifiers, which express additional factual knowledge. The hyper-relational knowledge graph completion (HKGC) task aims at inferring plausible missing links in a HKG. Most existing approaches to HKGC focus on enhancing the communication between qualifier pairs and main triples, while overlooking two important properties that emerge from the monotonicity of the hyper-relational graphs representation regime. Stage Reasoning allows for a two-step reasoning process, facilitating the integration of coarse-grained inference results derived solely from main triples and fine-grained inference results obtained from hyper-relational facts with qualifiers. In the initial stage, coarse-grained results provide an upper bound for correct predictions, which are subsequently refined in the fine-grained step. More generally, Qualifier Monotonicity implies that by attaching more qualifier pairs to a main triple, we may only narrow down the answer set, but never enlarge it. This paper proposes the HyperMono model for hyper-relational knowledge graph completion, which considers an initial triple-level reasoning phase to be subsequently refined by a triple+qualifiers reasoning step to infer missing entities. To realize qualifier monotonicity, HyperMono represents main triples as cones and each qualifier is modeled as a cone that shrinks the cone of the main triple to which the qualifier has been added. Moreover, HyperMono leverages the neighborhood context of an entity to semantically strengthen its representation. Experiments on three real-world datasets with three different scenario conditions demonstrate the strong performance of HyperMono when compared to the SoTA.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Data mining
@@ 1. Introduction
Knowledge graphs (KGs) store graph-like knowledge as a collection of factual triples (Bordes et al., 2013; Sun et al., 2019; Li et al., 2022). Each triple $(h,r,t)$ in a KG represents a connection between the head entity h and the tail entity t through the relation type $r$, e.g., (James Harden, member_of_team, Philadelphia 76ers) in Figure 1. Knowledge graphs provide the backbone of various applications like query answering (Zhang et al., 2021; Hu et al., 2022; Nguyen et al., 2023b), search (Nguyen et al., 2019; Gu et al., 2019), and recommendation systems (Yang et al., 2022; Zou et al., 2022). In hyper-relational knowledge graphs (HKGs), hyper-relational facts enhance binary relational facts by augmenting triples with supplementary knowledge expressed in the form of attribute-value qualifier pairs $(a:v)$. For example, in Figure 1, the triple (James Harden, member_of_team, Philadelphia 76ers) is contextualized by the qualifier pairs {(start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}, describing teams of the Atlantic Division in which James Harden and P.J. Tucker played together from 2019 to 2023. Like standard KGs, HKGs also suffer from incompleteness. The hyper-relational knowledge graph completion (HKGC) task, which aims at inferring missing links in HKGs, has therefore become an important research topic.
@@
    Qualifier Monotonicity generally implies that for a given hyper-relational query $q$, as the number of qualifier pairs in $q$ increases, the answer set of $q$ over a HKG might shrink but it never expands, while conversely removing qualifiers from $q$ could only lead to more possible answers. For example, for the query (James Harden, member_of_team, ?) in Figure 1, ? has five possible answers {Oklahoma City Thunder, Houston Rockets, Brooklyn Nets, Philadelphia 76ers, Los Angeles Clippers}. If the qualifier pairs (start_time: 2019) and (end_time: 2023) are added, then Oklahoma City Thunder is not a possible answer anymore. The main challenge for qualifier monotonicity lies in devising a representation that accurately captures its essence. It should posses the capability to faithfully capture the shrinking-like dynamics induced by the incorporation of qualifiers. In the context of embedding-based methods, ShrinkE (Xiong et al., 2023) has explored qualifier monotonicity through box embeddings. However, the inherent limitations of the embedding-based methods lead to a large gap between ShrinkE’s accuracy and that of SoTA models, as shown in § 4.1, Table 3.
@@ 2. Related Work
Transformer-based without GNNs. Many methods abandon the use of GNNs operations and instead incorporate hyper-relational knowledge through various transformer structures. For example, GRAN (Wang et al., 2021) implements edge-biased fully-connected attention into transformers to encode the graph structure and its heterogeneity. HyNT (Chung et al., 2023) defines a specialized context transformer to exchange information between the main triple and its qualifiers, HAHE (Luo et al., 2023) proposes a hierarchical attention structure with global-level and local-level attention to simultaneously model graph-like and sequential content. HyperFormer (Hu et al., 2023) introduces a mixture-of-experts strategy into the feed-forward layers of a transformer to strengthen its representation capabilities. However, these methods have two shortcomings: on the one hand, they only consider a fine-grained inference steps with hyper-relational knowledge and overlook coarse-grained inferences solely based on the main triples. On the other hand, they do not explicitly embody qualifier monotonicity. We note in passing that although the embeeding-based method ShrinkE (Xiong et al., 2023) takes qualifier monotonicity into account, its embedding-based nature results in a large gap in performance when compared with the state-of-the-art, as shown in § 4.1, Table 3.
@@ 3.1. Preliminaries
A hyper-relational knowledge graph (HKG) $\mathcal{G}$ is defined as {$\{\mathcal{E},\mathcal{R},\mathcal{H}\}$, where $\mathcal{E}$ is a set of entities, $\mathcal{R}$ is a set of relation types, and =$\mathcal{H}=\{(h,r,t,\mathcal{Q})~{}|~{}h\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{E}\}$ is a set of hyper-relational facts, such that $(h,r,t)$ denotes the main triple and $\mathcal{Q}=\{(a_{1}:v_{1}),\ldots,(a_{n}:v_{n})~{}|~{}a_{i}\in\mathcal{R},v_{i}\in\mathcal{E}\}$ is a set of auxiliary attribute-value qualifier pairs. We use $\mathcal{G}_{\textit{triple}}$ to denote the set of main triples occurring in $\mathcal{G}$. Note that when $n=0$, $\mathcal{G}$ is a standard knowledge graph, i.e., only containing main triples. Under this representation schema, the main triple (James Harden, member_of_team, Philadelphia 76ers) in Figure  1 with its qualifiers can be represented as: {(James Harden: member_of_team, Philadelphia 76ers), (start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}. Following the definition of the knowledge graph completion task, the hyper-relational knowledge graph completion (HKGC) task on $\mathcal{G}$ looks at predicting a missing entity ? in one of the following forms: (?$,r,t,\mathcal{Q})$ or $(h,r,\texttt{?},\mathcal{Q})$. Without loss of generality, in what follows we concentrate on the form of $(h,r,\texttt{?},\mathcal{Q}$).
@@
We present here only necessary background on cone embeddings, for a full discussion we refer the interested reader to (Zhang et al., 2021; Nguyen et al., 2023c, a). A d-dimensional sector-cone can be parameterized by $\mathcal{C}^{d}(\alpha,\beta)=((\alpha^{1},\beta^{1}),...,(\alpha^{d},\beta^{d}))$, where $\alpha^{i}\in[-\pi,\pi)$ represents the angle between the symmetry axis of the i-th-dimension sector-cone and the positive x axis, and $\beta^{i}$ represents the aperture of the i-th-dimension sector-cone. We omit the superscript of $\mathcal{C}^{d}$ if it is clear from context. A variety of operations on cone embeddings can be defined, here we introduce the projection $\mathcal{P}$ and intersection $\mathcal{I}$ operations:
@@ 3.2. Model Architecture
We present in this section the two main components of HyperMono’s architecture: Head Neighborhood Encoder (§ 3.2.1) and Missing Entity Predictor (§ 3.2.2).
@@ 3.2.1. HNE: Head Neighborhood Encoder
To adequately encode the content of the neighbors of a head entity, as a first step, we introduce the Head Neighborhood Encoder (HNE) module, cf.  Figure 2(b). The HNE module considers two types of relational contexts, characterized by the use of either main triples or complete hyper-relational facts to ascertain the neighbors of a given entity: Coarse-grained Neighbor Aggregator (CNA), only considering main triples in $\mathcal{G}_{triple}$; Fine-grained Neighbor Aggregator (FNA), considering hyper-relational facts in $\mathcal{G}$. The resulting neighbor-aware embeddings from the CNA and FNA modules will be used in the subsequent module (§3.2.2) to provide further evidence, based on the neighbors of $h$, for the prediction of the ???? entity.
@@
    Local and Global Entity Inference. Different neighbors of $h$ might provide information about $h$ from different perspectives. We thus propose the Local Entity Inference (LEI) mechanism that independently considers each neighbor to infer the entity corresponding to h. LEI focuses on a single neighbor during inference, reducing the interference from irrelevant information. In practice, for $\textit{tr}\in N_{h}$, we perform a matrix multiplication operation on $\textbf{{N}}_{\textit{tr}}^{h}$ and $\textbf{{V}}_{\!\textit{ent}}$ to obtain the entity score vector $\textbf{{P}}_{\textit{tr}}^{h}\in\mathbb{R}^{1\times N}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{{N}}_{\textit{tr}}^{h}}\cdot\textbf{{V}}_{\!\textit{ent}}^{\top}$, where $\textbf{{V}}_{\!\textit{ent}}\in\mathbb{R}^{N\times d}$ denotes the embedding matrix for all entities in the dataset, $\textbf{{V}}_{\!\textit{ent}}^{\top}$ is the transpose of $\textbf{{V}}_{\!\textit{ent}}$, and $N=|\mathcal{E}|$. However, in some cases, it is difficult to infer a complex entity from a single neighbor. Therefore, we additionally propose a Global Entity Inference (GEI) mechanism which averages the entity score vector predicted by different relational neighbors to aggregate them: $\textbf{{P}}_{G}^{h}=mean(\sum_{\textit{tr}\in N_{h}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{{P}}_{\textit{tr}}^{h}})$.
@@ 3.2.2. MEP: Missing Entity Predictor
    (3) $\left\{\begin{array}[]{l}\alpha_{r,a_{i},v_{i}}=f(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}})),\,\,\beta_{r,a_{i},v_{i}}=g(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}}))\vspace{1.2ex}\\ {\rm offset}=(\beta_{h,r}-\beta_{\mathcal{S}_{i}})\odot\sigma(\alpha_{r,a_{i},v_{i}}),\,\,\beta_{\mathcal{S}_{i}}=\sigma(\beta_{r,a_{i},v_{i}})\vspace{1.2ex}\\ \alpha_{\mathcal{S}_{i}}=\alpha_{h,r}-0.5\odot\beta_{h,r}+0.5\odot\beta_{\mathcal{S}_{i}}+0.5\odot{\rm offset}\\ \end{array}\right.$
    where $\Theta_{r,a_{i},v_{i}}=\texttt{MLP}_{2}(\texttt{Concat}(\textbf{{E}}^{r},\textbf{{E}}^{a_{i}},\textbf{{E}}^{v_{i}}))$, which leverages an MLP layer which takes the primary relation and attribute-value qualifier as input, and outputs the combined representation of relational and hyper-relational knowledge. $\texttt{MLP}_{1}$: $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\texttt{MLP}_{2}$: $\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ are multi-layer perceptron networks, $\sigma$ is a sigmoid function and $\odot$ is the element-wise vector multiplication.
    From Equation (3), we obtain a sector-cone $\mathcal{C}_{\mathcal{S}_{i}}=\mathcal{C}(\alpha_{\mathcal{S}_{i}},\beta_{\mathcal{S}_{i}})$ that has been shrinked by the qualifier ($a_{i}$ : $v_{i}$).
@@
    After obtaining a scaled down sector-cone $\mathcal{C}_{\mathcal{S}_{i}}$, for each $i$-th qualifier, we obtain $s$ independent cones: $\mathcal{C}_{\mathcal{S}_{1}},...,\mathcal{C}_{\mathcal{S}_{s}}$}. Finally, when considering multiple qualifier pairs at the same time, the answer set can be computed by the cone intersection operation $\mathcal{I}$: $\mathcal{C}_{\mathcal{S}}=(\alpha_{\mathcal{S}},\beta_{\mathcal{S}})=\mathcal{I}(\mathcal{C}_{\mathcal{S}_{1}},...,\mathcal{C}_{\mathcal{S}_{s}})$.
Inference Results. We apply $\textbf{{q}}^{t}=\theta(\texttt{MLP}(\texttt{Concat}(\alpha_{\mathcal{S}},\beta_{\mathcal{S}}))\cdot\textbf{{V}}_{ent}^{\top})$ to change the answer cone $\mathcal{C}_{\mathcal{S}}$ to the entity probability at the ? position. Additionally, we calculate the cross-entropy loss between the prediction score $\textbf{{q}}^{t}$ and the label $\textbf{{L}}^{t}_{2}$ as $\mathcal{L}_{hyper}(t)=\texttt{CrossEntropy}(\textbf{{q}}^{t},\textbf{{L}}^{t}_{2})$.

### v3

@@ Abstract.
In a hyper-relational knowledge graph (HKG), each fact is composed of a main triple associated with attribute-value qualifiers, which express additional factual knowledge. The hyper-relational knowledge graph completion (HKGC) task aims at inferring plausible missing links in a HKG. Most existing approaches to HKGC focus on enhancing the communication between qualifier pairs and main triples, while overlooking two important properties that emerge from the monotonicity of the hyper-relational graphs representation regime. Stage Reasoning allows for a two-step reasoning process, facilitating the integration of coarse-grained inference results derived solely from main triples and fine-grained inference results obtained from hyper-relational facts with qualifiers. In the initial stage, coarse-grained results provide an upper bound for correct predictions, which are subsequently refined in the fine-grained step. More generally, Qualifier Monotonicity implies that by attaching more qualifier pairs to a main triple, may only narrow down the answer set, but never enlarges it. This paper proposes the HyperMono model for hyper-relational knowledge graph completion, which considers an initial triple-level reasoning phase to be subsequently refined by a triple+qualifiers reasoning step to infer missing entities. To realize qualifier monotonicity, HyperMono represents main triples as cones and each qualifier is modeled as a cone that shrinks the cone of the main triple to which the qualifier has been added. Moreover, HyperMono leverages the neighborhood context of an entity to semantically strengthen its representation. Experiments on three real-world datasets with three different scenario conditions demonstrate the strong performance of HyperMono when compared to the SoTA.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Knowledge graphs (KGs) store graph-like knowledge as a collection of factual triples (Bordes et al., 2013; Sun et al., 2019). Each triple $(h,r,t)$ in a KG represents a connection between the head entity h and the tail entity t through the relation type $r$, e.g., (James Harden, member_of_team, Philadelphia 76ers) in Figure 1.Knowledge graphs provide the backbone of various applications like query answering (Zhang et al., 2021; Hu et al., 2022), search (Nguyen et al., 2019), and recommendation systems (Yang et al., 2022a). In hyper-relational knowledge graphs (HKGs), hyper-relational facts enhance binary relational facts by augmenting triples with supplementary knowledge expressed in the form of attribute-value qualifier pairs $(a:v)$. For example, in Figure 1, the triple (James Harden, member_of_team, Philadelphia 76ers) is contextualized by the qualifier pairs {(start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}, describing teams of the Atlantic Division in which James Harden and P.J. Tucker played together from 2019 to 2023. Like standard KGs, HKGs also suffer from incompleteness. The hyper-relational knowledge graph completion (HKGC) task, which aims at inferring missing links in HKGs, has therefore become an important research topic.
@@
    Qualifier Monotonicity generally implies that for a given hyper-relational query $q$, as the number of qualifier pairs in $q$ increases, the answer set of $q$ over a HKG might shrink but it never expands, while conversely removing qualifiers from $q$ could only lead to more possible answers. For example, for the query (James Harden, member_of_team, ?) in Figure 1, ? has five possible answers {Oklahoma City Thunder, Houston Rockets, Brooklyn Nets, Philadelphia 76ers, Los Angeles Clippers}. If the qualifier pairs (start_time: 2019) and (end_time: 2023) are added, then Oklahoma City Thunder is not a possible answer anymore. The main challenge for qualifier monotonicity lies in devising a representation that accurately captures its essence. It should posses the capability to faithfully capture the shrinking-like dynamics induced by the incorporation of qualifiers. In the context of embedding-based methods, ShrinkE (Xiong et al., 2023) has explored qualifier monotonicity through box embeddings. However, the inherent limitations of the embedding-based methods lead to a large gap between ShrinkE’s accuracy and that of SoTA models, as shown in Table 1 of § 4.2.
@@ 2. Related Work
Transformer-based without GNNs. Many methods abandon the use of GNNs operations and instead incorporate hyper-relational knowledge through various transformer structures. For example, GRAN (Wang et al., 2021) implements edge-biased fully-connected attention into transformers to encode the graph structure and its heterogeneity. HyNT (Chung et al., 2023) defines a specialized context transformer to exchange information between the main triple and its qualifiers, HAHE (Luo et al., 2023) proposes a hierarchical attention structure with global-level and local-level attention to simultaneously model graph-like and sequential content. HyperFormer (Hu et al., 2023) introduces a mixture-of-experts strategy into the feed-forward layers of a transformer to strengthen its representation capabilities. However, these methods have two shortcomings: on the one hand, they only consider a fine-grained inference steps with hyper-relational knowledge and overlook coarse-grained inferences solely based on the main triples. On the other hand, they do not explicitly embody qualifier monotonicity. We note in passing that although the embeeding-based method ShrinkE (Xiong et al., 2023) takes qualifier monotonicity into account, its embedding-based nature results in a large gap in performance when compared with the state-of-the-art, as shown in Table 1 of § 4.2.
@@ 3.1. Preliminaries
A hyper-relational knowledge graph (HKG) $\mathcal{G}$ is defined as {$\{\mathcal{E},\mathcal{R},\mathcal{H}\}$, where $\mathcal{E}$ is a set of entities, $\mathcal{R}$ is a set of relation types, and =$\mathcal{H}=\{(h,r,t,\mathcal{Q})~{}|~{}h\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{E}\}$ is a set of hyper-relational facts, such that $(h,r,t)$ denotes the main triple and $\mathcal{Q}=\{(a_{1}:v_{1}),\ldots,(a_{n}:v_{n})~{}|~{}a_{i}\in\mathcal{R},v_{i}\in\mathcal{E}\}$ is a set of auxiliary attribute-value qualifier pairs. We use $\mathcal{G}_{\textit{triple}}$ to denote the set of main triples occurring in $\mathcal{G}$. Note that when $n=0$, $\mathcal{G}$ is a standard knowledge graph, i.e., only containing main triples. Under this representation schema, the main triple (James Harden, member_of_team, Philadelphia 76ers) in Figure  1 with its qualifiers can be represented as: {(James Harden: member_of_team, Philadelphia 76ers), (start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}. Building upon the definition of the knowledge graph completion task (Bordes et al., 2013; Sun et al., 2019), the hyper-relational knowledge graph completion (HKGC) task aims at predicting entities within hyper-relational facts $\mathcal{H}$. Based on the entities to be predicted, HKGC can be classified into two distinct categories:
  1. (1)
    Head / Tail: the entities to be predicted are exclusively located within the main triple, i.e., predict a missing entity ? in (?$,r,t,\mathcal{Q})$ or $(h,r,\texttt{?},\mathcal{Q})$.
  2. (2)
    All Entities: the entities to be predicted may not only reside within the main triple but could also be present in the qualifier pairs $\mathcal{Q}$, i.e., predict a missing entity ? in (?$,r,t,\mathcal{Q})$ or $(h,r,\texttt{?},\mathcal{Q})$ or $\{(a_{1}:v_{1}),\ldots,(a_{i}:\texttt{?}),\ldots,(a_{n}:v_{n})\}$.
Without loss of generality, in what follows we concentrate on predicting a tail entity ? in $(h,r,\texttt{?},\mathcal{Q})$. The other settings can be handled in a similar manner.
@@
We present here only necessary background on cone embeddings, for a full discussion we refer the interested reader to (Zhang et al., 2021; Nguyen et al., 2023b, a). A d-dimensional sector-cone can be parameterized by $\mathcal{C}^{d}(\alpha,\beta)=((\alpha^{1},\beta^{1}),...,(\alpha^{d},\beta^{d}))$, where $\alpha^{i}\in[-\pi,\pi)$ represents the angle between the symmetry axis of the i-th-dimension sector-cone and the positive x axis, and $\beta^{i}\in[0,2\pi)$ represents the aperture of the i-th-dimension sector-cone. We omit the superscript of $\mathcal{C}^{d}$ if it is clear from context. A variety of operations on cone embeddings can be defined, here we introduce the projection $\mathcal{P}$ and intersection $\mathcal{I}$ operations:
@@ 3.2. Model Architecture
We present the two main components of HyperMono’s architecture: Head Neighborhood Encoder (§ 3.2.1) and Missing Entity Predictor (§ 3.2.2).
@@ 3.2.1. HNE: Head Neighborhood Encoder
To adequately encode the content of the neighbors of a head entity, as a first step, we introduce the Head Neighborhood Encoder (HNE) module, cf.  Figure 2(b). The HNE module considers two types of relational contexts, characterized by the use of either main triples or complete hyper-relational facts to ascertain the neighbors of a given entity: Coarse-grained Neighbor Aggregator (CNA), only considering main triples in $\mathcal{G}_{triple}$; Fine-grained Neighbor Aggregator (FNA), considering hyper-relational facts in $\mathcal{G}$. The resulting neighbor-aware embeddings from the CNA and FNA modules will be used in the subsequent module (§3.2.2) to provide further evidence, based on the neighbors of $h$, for the prediction of the ? entity.
@@
    Local and Global Entity Inference. Different neighbors of $h$ might provide information about $h$ from different perspectives. We thus propose the Local Entity Inference (LEI) mechanism that independently considers each neighbor to infer the entity corresponding to h. LEI focuses on a single neighbor during inference, reducing the interference from irrelevant information. In practice, for $\textit{tr}\in N_{h}$, we perform a matrix multiplication operation on $\textbf{{N}}_{\textit{tr}}^{h}$ and $\textbf{{V}}_{\!\textit{ent}}$ to obtain the entity score vector $\textbf{{P}}_{\textit{tr}}^{h}\in\mathbb{R}^{1\times N}=\textbf{{N}}_{\textit{tr}}^{h}\cdot\textbf{{V}}_{\!\textit{ent}}^{\top}$, where $\textbf{{V}}_{\!\textit{ent}}\in\mathbb{R}^{N\times d}$ denotes the embedding matrix for all entities in the dataset, $\textbf{{V}}_{\!\textit{ent}}^{\top}$ is the transpose of $\textbf{{V}}_{\!\textit{ent}}$, and $N=|\mathcal{E}|$. However, in some cases, it is difficult to infer a complex entity from a single neighbor. Therefore, we additionally propose a Global Entity Inference (GEI) mechanism which averages the entity score vector predicted by different relational neighbors to aggregate them: $\textbf{{P}}_{G}^{h}=mean(\sum_{\textit{tr}\in N_{h}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{{P}}_{\textit{tr}}^{h}})$.
@@ 3.2.2. MEP: Missing Entity Predictor
    (3) $\left\{\begin{array}[]{l}\alpha_{r,a_{i},v_{i}}=f(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}})),\,\,\beta_{r,a_{i},v_{i}}=g(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}}))\vspace{1.2ex}\\ {\rm offset}=(\beta_{h,r}-\beta_{\mathcal{S}_{i}})\odot\sigma(\alpha_{r,a_{i},v_{i}}),\,\,\beta_{\mathcal{S}_{i}}=\sigma(\beta_{r,a_{i},v_{i}})\odot\beta_{h,r}\vspace{1.2ex}\\ \alpha_{\mathcal{S}_{i}}=\alpha_{h,r}-0.5\odot\beta_{h,r}+0.5\odot\beta_{\mathcal{S}_{i}}+0.5\odot{\rm offset}\\ \end{array}\right.$
    where $\Theta_{r,a_{i},v_{i}}=\texttt{MLP}_{2}(\texttt{Concat}(\textbf{{E}}^{r},\textbf{{E}}^{a_{i}},\textbf{{E}}^{v_{i}}))$, which leverages an MLP layer which takes the primary relation and attribute-value qualifier as input, and outputs the combined representation of relational and hyper-relational knowledge. $\texttt{MLP}_{1}$: $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\texttt{MLP}_{2}$: $\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ are multi-layer perceptron networks, $\sigma$ is a sigmoid function and $\odot$ is the element-wise vector multiplication. From Equation (3), we obtain a sector-cone $\mathcal{C}_{\mathcal{S}_{i}}=\mathcal{C}(\alpha_{\mathcal{S}_{i}},\beta_{\mathcal{S}_{i}})$ that has been shrinked by the qualifier ($a_{i}$ : $v_{i}$).
@@
    After obtaining a scaled down sector-cone $\mathcal{C}_{\mathcal{S}_{i}}$, for each $i$-th qualifier, we obtain $s$ independent cones: $\mathcal{C}_{\mathcal{S}_{1}},...,\mathcal{C}_{\mathcal{S}_{s}}$}. Finally, when considering multiple qualifier pairs at the same time, the answer set can be computed by the cone intersection operation $\mathcal{I}$: $\mathcal{C}_{\mathcal{S}}=(\alpha_{\mathcal{S}},\beta_{\mathcal{S}})=\mathcal{I}(\mathcal{C}_{\mathcal{S}_{1}},...,\mathcal{C}_{\mathcal{S}_{s}})$. Note that when $C_{S}$ is empty, the intersection of all $C_{S_{i}}$ must form a cone with an aperture of 0. However, this scenario is highly improbable. The key reason is that each cone exists in a high-dimensional space (with a dimensionality of 200 for our HyperMono), making it virtually impossible for all $\beta$ values in $C_{S}$ to be 0. Even in extreme cases where the aperture is 0, he cone embedding remains effective in handling this situation, as discussed in ConE (Zhang et al., 2021).
Inference Results. We apply $\textit{q}^{t}=f(\texttt{MLP}(\texttt{Concat}(\alpha_{\mathcal{S}},\beta_{\mathcal{S}})),\textbf{{V}}_{ent}^{\top})$, where $f(\cdot)=\mathcal{D}(\textbf{{V}}_{ent},\mathcal{C}_{\mathcal{S}})$ denotes the score function to calculate the distance between $\textbf{{V}}_{ent}$ and the scaled cone $\mathcal{C}_{\mathcal{S}}$, the distance function is consistent with ConE (Zhang et al., 2021), $\textbf{{q}}^{t}=\theta(\textit{q}^{t})$ to change the answer cone $\mathcal{C}_{\mathcal{S}}$ to the entity probability at the ? position, $\theta$ is the softmax function. Additionally, we calculate the cross-entropy loss between the prediction score $\textbf{{q}}^{t}$ and the label $\textbf{{L}}^{t}_{2}$ as $\mathcal{L}_{hyper}(t)=\texttt{CrossEntropy}(\textbf{{q}}^{t},\textbf{{L}}^{t}_{2})$.
Table 1. Evaluation of different models under the mixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. All baseline results are collected from the original papers. Best scores are highlighted in bold, the second best scores are highlighted in normal, the $\Delta$ represents the performance improvement over the state-of-the-art modes, and ’–’ indicates the results are not reported in previous work. The number in parentheses represents the proportion of triples with hyper-relational knowledge.
Methods WD50K (13.6) WikiPeople (2.6) JF17K (45.9)
Head / Tail All Entities Head / Tail All Entities Head / Tail All Entities
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
m-TransH (Wen et al., 2016) 0.063 0.063 0.300 0.206 0.206 0.462 0.102 0.069 0.168
RAE (Zhang et al., 2018) 0.058 0.058 0.306 0.172 0.102 0.320 0.215 0.215 0.466 0.310 0.219 0.504
NaLP-Fix (Rosso et al., 2020) 0.177 0.131 0.264 0.224 0.158 0.330 0.408 0.331 0.546 0.338 0.272 0.466 0.221 0.165 0.331 0.366 0.290 0.516
NeuInfer (Guan et al., 2020) 0.243 0.176 0.377 0.228 0.162 0.341 0.476 0.415 0.585 0.333 0.259 0.477 0.449 0.361 0.624 0.473 0.397 0.618
HINGE (Rosso et al., 2020) 0.243 0.176 0.377 0.232 0.164 0.343 0.342 0.272 0.463 0.350 0.282 0.467 0.431 0.342 0.611 0.517 0.436 0.675
StarE (Galkin et al., 2020) 0.349 0.271 0.496 0.491 0.398 0.592 0.378 0.265 0.542 0.574 0.496 0.725 0.542 0.454 0.685
GRAN (Wang et al., 2021) 0.309 0.240 0.441 0.503 0.438 0.620 0.479 0.410 0.604 0.617 0.539 0.770 0.656 0.582 0.799
HAHE (Luo et al., 2023) 0.368 0.291 0.516 0.402 0.327 0.546 0.509 0.447 0.639 0.495 0.420 0.631 0.623 0.554 0.806 0.668 0.597 0.816
HyperFormer (Hu et al., 2023) 0.366 0.288 0.514 0.417 0.344 0.555 0.491 0.386 0.654 0.479 0.373 0.644 0.664 0.601 0.787 0.662 0.596 0.790
HyperCL (Lu et al., 2024) 0.302 0.231 0.441 0.395 0.321 0.539 0.455 0.365 0.596 0.509 0.437 0.644 0.513 0.429 0.684 0.673 0.604 0.818
HyperMono 0.375 0.298 0.522 0.447 0.376 0.582 0.494 0.390 0.657 0.498 0.394 0.659 0.685 0.630 0.796 0.766 0.717 0.861
$\Delta$ (%) +0.7 +0.7 +0.6 +3.0 +3.2 +2.7 -1.5 -5.7 +0.3 -1.1 -4.3 +1.5 +2.1 +2.9 -1.0 +9.3 +11.3 +4.3